# November, 2025

### v1

@@ 5 Black hole molecule
It is now natural to examine the “black hole molecule” that consists of two black hole hydrogen atoms. This is composed of the two stable charged black holes together with the quantum electrons that are trapped by them. We consider the multi-centered Majumdar–Papapetrou black hole solution as the nuclei of the molecule. The Majumdar–Papapetrou solution in four dimensions is given by
@@
where $U$ is the harmonic function in $\mathbb{R}^{3}$:
@@
Here $M_{i}$ ($i=1,\ldots,N$) are constants and $(x_{i},y_{i},z_{i})$ are centers of the black holes in $\mathbb{R}^{3}$. Each black hole in the Majumdar–Papapetrou solution (5.1) is stable since the gravitational attraction and the electromagnetic repulsion are balanced.
@@
where $\vec{R}_{i}=(x-x_{i},y-y_{i},z-z_{i})$, ($i=1,2$). As in the case of the ordinary hydrogen molecule in the flat space, we employ the Born–Oppenheimer approximation. Namely, we ignore the dynamics of the nuclei (the binary black holes). We also consider a single electron trapped on them for simplicity. This corresponds to the hydrogen molecular ion $\mathrm{H}_{2}^{+}$ composed of the black holes having each charge +e+e and a single electron with mass $\mu$ and the charge e-e. Then the Schrödinger equation is given by
@@
where we have included the potential $V(r)=-\frac{e^{2}}{R_{1}}-\frac{e^{2}}{R_{2}}$ that provides the electromagnetic attractive force [^4] Usually, the potential includes $V(r)=\frac{e^{2}}{R_{12}}$ that represents the repulsive force between two nuclei. It is obvious that in our setup, such repulsive force does not exist. . For the standard $\mathrm{H}_{2}^{+}$ case, a variable separation is possible by using the spheroidal coordinates. We assume that the two black holes are located on the points (0,0,d/2)($0,$-d/2) and $(0,0,d/2)$ in $\mathbb{R}^{3}$ and $M_{1}=M_{2}=M$ (Fig. 11).
@@
It is convenient to employ the coordinates,
@@
These together with an azimuth angle $\phi$ consists of the spheroidal coordinates. They are defined in $1\leq\xi<\infty$, $1\leq\eta\leq 1$, $0\leq\phi<2\pi$. These are related to the Cartesian coordinate as
@@
where we have defined $A=\frac{2\mu e^{2}}{\hbar^{2}}$, $\kappa^{2}=-\frac{2\mu E}{\hbar^{2}}$ and used the following relation:
@@
We find that the $\phi$ dependence is decoupled but the $\xi$ and $\eta$ sectors are not. This is expected since the Stäckel condition of the equation is violated by the curvature corrections. Therefore the equation is solved by the numerical analysis.
In order to solve the equation analytically, it is necessary to decouple the two variables $\xi$ and $\eta$. To this end, we consider a specific limit $\mu\to 0$ with fixed $M$. This corresponds to the limit where the mass of the electron is small compared to the mass of the black hole, and is a natural assumption. Then the third line in (LABEL:eq:molecule) is ignored and the terms in the second line in the large-$\xi$ region becomes $\frac{M}{d}\frac{1}{R}\frac{\mathrm{d}R}{\mathrm{d}\xi}$. Then the equation in this asymptotic region becomes
@@
$\frac{\mathrm{d}}{\mathrm{d}\eta}\Big\{(1-\eta^{2})\frac{\mathrm{d}}{\mathrm{d}\eta}S\Big\}+\Big(\Lambda-\frac{\alpha^{2}}{1-\eta^{2}}\Big)S=0.$ (5.11)
where $\Lambda$ is a parameter. Now in this specific regime and parameters, the two equations are decoupled and can be solved. It is easy to show that the equation for $R(\xi)$ is solved by an integration of the hypergeometric functions $_{2}F_{1}(a,b,c,\xi)$ for some constants $a,b,c$. The equation for $S(\eta)$ is solved by the associated Legendre polynomials $P^{m}_{n}(\eta)$ and the associated Legendre function $Q^{m}_{n}(\eta)$ of the second kind.
@@ 6 Conclusion and discussions
We first derived the equation for free particle in the spatial geometry of the Schwarzschild black hole. We found that the equation is rewritten in the form of the confluent Heun equation and the it is therefore solved analytically. In order to visualize the wavefunction, we found it convenient to solve the equation numerically. We showed that the wavefunction is attracted toward the black hole and observed that particles with higher energy are more strongly attracted. We then introduced a hypothetical electric attractive potential to the Schwarzschild black hole and studied the bound states. The equation is again rewritten in the form of the confluent Heun equation. We numerically find solutions satisfying appropriate boundary condition and visualized the wavefunctions for the hydrogen atom type configuration. We found that the wavefunction and the electron cloud is attracted into the horizon and it is confined around the horizon.
@@
Next, we focused on the charged black hole and discussed the genuine black hole atom. We considered the Reissner–Nordström black hole solution and derived the Schrödinger equation. Assuming a spherically symmetric solution, we numerically solved the equation in the radial direction. We then examined the effects of the black hole on the hydrogen atom for various quantum numbers. We visualized the wavefunction and explicitly showing the difference from the ordinary hydrogen atom. We observed that the electron cloud is generically bounded to the outer horizon. We also investigated the wavefunction extending into the region inside the inner horizon. We observed that the wavefunction is attracted to the inner horizon from the interior. This is most naturally interpreted as the famous gravitational repulsion force present in $r\sim 0$ region of the Reissner–Nordström black hole. These results indicate that not only classical objects but also quantum wavefunctions are affected by gravity and fall into black holes. This is obvious intuitively (classically), but not evident in quantum theory. We also studied the wavefunctions in the extremal and the over-extremal cases.
@@
We then derived the Schrödinger equation for a molecule with a binary black hole as its nucleus. Since there is no force acting between these black holes, it is in some sense simpler than ordinary molecules. However, due to the presence of the curvature term, the spheroidal coordinates variables can not be separated. Therefore, numerical analysis is generally required. We demonstrated that under specific conditions, the equation becomes separable and analytical solutions can be obtained.
@@
It would be helpful whether the black holes discussed in this paper is available in the current Universe or not. The Schwarzschild radius of a black hole of mass $M$ is given by $r_{0}=\frac{2G_{\mathrm{N}}M}{c^{2}}$ where $G_{\mathrm{N}}$ is the Newton’s gravitational constant. On the other hand, the Bohr radius of the hydrogen atom is known as $a_{0}=\frac{4\pi\varepsilon_{0}\hbar^{2}}{m_{e}e^{2}}\sim 10^{-10}\,\mathrm{m}$ (where $m_{e}$ is the mass of the electron). The mass of the black holes of the size of the hydrogen atom is estimated as $M=\frac{c^{2}a_{0}}{2G_{\mathrm{N}}}\sim 6.7\times 10^{16}\,\mathrm{kg}$ which is the same order of the small asteroid. Black holes of such a small mass are not created by the gravitational collapse of stars but may appear in the early stage of the Universe. The time $\Delta t$ that it takes for a black hole of mass $M\sim 10^{16}\,\mathrm{kg}$ to disappear through the Hawking radiation is estimated as $\Delta t\sim\frac{G_{\mathrm{N}}}{\hbar c^{3}}M^{3}\sim 10^{24}\,\mathrm{yr}$, which is greater than the age of the Universe. Therefore it is possible that it remains in the current Universe as a primordial black hole. It can not be excluded that such black holes trap electrons and forms the “atoms” discussed in this paper.
@@
As an interesting consideration related to what we discussed here is the discrepancy between the classical and quantum depictions of black hole atoms. In the classical picture, electrons can not remain stationary near a black hole and they must fall into it. On the other hand, as we see in the ordinary hydrogen atom, quantum effects do not allow electrons to fall into the interior of the nucleus, but keeps them in stable orbits. These conflicting facts can coexist in the black hole atom. That is, while the electron cloud maintains a stable orbit due to quantum effects, parts of it are attracted to the horizon and fall into the interior of black holes. This “gravity versus quantum” effects is interesting issue to study. It would be interesting to study this issue in more complex black holes, such as Kerr solutions and so on.
@@ Acknowledgments
The work of S. S. and K. S. is supported by Grant-in-Aid for Scientific Research, JSPS KAKENHI Grant Number JP25K07324. The work of K. S. is also supported by Kitasato University Research Grant for Young Researchers.

### v2

@@ 5 Black hole molecule
According to elementary chemistry, when two atomic nuclei share electrons, a hydrogen molecule is formed. In particular, the hydrogen molecular ion $\mathrm{H}_{2}^{+}$, where two atomic nuclei share a single electron, is known as a model that can be solved exactly. It is now natural to examine the “black hole molecule” that consists of two black hole hydrogen atoms. This is composed of the two stable charged black holes together with the quantum electrons that are trapped by them. We consider the multicentered Majumdar–Papapetrou black hole solution as the nuclei of the molecule. The Majumdar–Papapetrou solution in four dimensions is given by
@@
where $U$ is the harmonic function in $\mathbb{R}^{3}$,
@@
Here, $M_{i}$ ($i=1,\ldots,N$) are constants and $(x_{i},y_{i},z_{i})$ are centers of the black holes in $\mathbb{R}^{3}$. Each black hole in the Majumdar–Papapetrou solution (5.1) is stable since the gravitational attraction and the electromagnetic repulsion are balanced.
@@
where $\vec{R}_{i}=(x-x_{i},y-y_{i},z-z_{i})$ ($i=1,2$). As in the case of the ordinary hydrogen molecule in the flat space, we employ the Born–Oppenheimer approximation. Namely, we ignore the dynamics of the nuclei (the binary black holes). We also consider a single electron trapped on them for simplicity. This corresponds to the hydrogen molecular ion $\mathrm{H}_{2}^{+}$ composed of the black holes having each charge +e+e and a single electron with mass $\mu$ and the charge e-e. Then the Schrödinger equation is given by
@@
where we have included the potential $V(r)=-\frac{e^{2}}{R_{1}}-\frac{e^{2}}{R_{2}}$ that provides the electromagnetic attractive force. [^4] For the ordinary hydrogen molecular ion $\mathrm{H}_{2}^{+}$, the potential includes $V(r)=\frac{e^{2}}{R_{12}}$ that represents the repulsive force between two nuclei. It is obvious that in our setup, such repulsive force does not exist. For the standard $\mathrm{H}_{2}^{+}$ case, a variable separation is possible by using the spheroidal coordinates. We assume that the two black holes are located on the points (0,0,d/2)($0,$-d/2) and $(0,0,d/2)$ in $\mathbb{R}^{3}$ and $M_{1}=M_{2}=M$ (Fig. 11).
@@
It is convenient to employ the coordinates
@@
These together with an azimuth angle $\phi$ consist of the spheroidal coordinates. They are defined in $1\leq\xi<\infty$, $1\leq\eta\leq 1$, and $0\leq\phi<2\pi$. These are related to the Cartesian coordinate as
@@
where we have defined $A=\frac{2\mu e^{2}}{\hbar^{2}}$, $\kappa^{2}=-\frac{2\mu E}{\hbar^{2}}$, and used the following relation:
@@
We find that the $\phi$ dependence is decoupled but the $\xi$ and $\eta$ sectors are not. This is expected since the Stäckel condition of the equation is violated by the curvature corrections. Therefore, the equation is solved by the numerical analysis.
In order to solve the equation analytically, it is necessary to decouple the two variables $\xi$ and $\eta$. To this end, we consider a specific limit $\mu\to 0$ with fixed $M$. This corresponds to the limit where the mass of the electron is small compared to the mass of the black hole, and is a natural assumption. Then the third line in (LABEL:eq:molecule) is ignored, and the terms in the second line in the large-$\xi$ region become $\frac{M}{d}\frac{1}{R}\frac{\mathrm{d}R}{\mathrm{d}\xi}$. Then the equation in this asymptotic region becomes
@@
$\frac{\mathrm{d}}{\mathrm{d}\eta}\Big\{(1-\eta^{2})\frac{\mathrm{d}}{\mathrm{d}\eta}S\Big\}+\Big(\Lambda-\frac{\alpha^{2}}{1-\eta^{2}}\Big)S=0,$ (5.11)
where $\Lambda$ is a parameter. Now in this specific regime and parameters, the two equations are decoupled and can be solved analytically. Note that the effects of the black holes $M\not=0$ still remain in these equations. We find solutions given by
$R(\xi)=$ $\ \left(\frac{1-\xi}{1+\xi}\right)^{\frac{M}{2d}}\Big(c_{1}P^{B}_{A}(\xi)+c_{2}Q^{B}_{A}(\xi)\Big),$
$S(\eta)=$ $\ c_{3}P^{\alpha}_{A}(\eta)+c_{4}Q^{\alpha}_{A}(\eta),$ (5.12)
where $P^{B}_{A}(x)$ and $Q^{B}_{A}(x)$ are the associated Legendre functions of the first and the second kinds, respectively. The parameters are given by $A=\frac{1}{2}(\sqrt{4\Lambda+1}-1)$ and $B=\frac{1}{2d}\sqrt{M^{2}+4d^{2}\alpha^{2}}$, and $c_{1},\ldots,c_{4}$ are integration constants. The single-valuedness of the wave function requires that $\alpha$ be an integer. If we require the discrete energy spectra and the normalizability for bound states, $A$ and $B$ would be integers satisfying the conditions $0\leq B\leq A$ and $0\leq\alpha\leq A$. It is known that the exact solution for the ordinary hydrogen molecule ion $\mathrm{H}_{2}^{+}$ is given by the associated Legendre polynomials $P^{m}_{n}(x)$ and the confluent Heun function $H(x)$ [28]. Although the solution (5.12) has a similar structure with the ordinary $\mathrm{H}_{2}^{+}$ in the sense that they share the Legendre polynomials, they are of course different due to the curvature corrections and the absence of the internuclei potential.
@@ 6 Conclusion and discussions
We first derived the equation for free particle in the spatial geometry of the Schwarzschild black hole. We found that the equation is rewritten in the form of the confluent Heun equation and it is, therefore, solved analytically. In order to visualize the wave function, we found it convenient to solve the equation numerically. We showed that the wave function is attracted toward the black hole and observed that particles with higher energy are more strongly attracted. We then introduced a hypothetical electric attractive potential to the Schwarzschild black hole and studied the bound states. The equation is again rewritten in the form of the confluent Heun equation. We numerically found solutions satisfying appropriate boundary conditions and visualized the wave functions for the hydrogen atom type configuration. We found that the wave function and the electron cloud are attracted into the horizon and are confined around the horizon.
@@
Next, we focused on the charged black hole and discussed the genuine black hole atom. We considered the Reissner–Nordström black hole solution and derived the Schrödinger equation. Assuming a spherically symmetric solution, we numerically solved the equation in the radial direction. We then examined the effects of the black hole on the hydrogen atom for various quantum numbers. We visualized the wave function, explicitly showing the difference from the ordinary hydrogen atom. We observed that the electron cloud is generically bound to the outer horizon. We also investigated the wave function extending into the region inside the inner horizon. We observed that the wave function is attracted to the inner horizon from the interior. This is most naturally interpreted as the famous gravitational repulsion force present in the $r\sim 0$ region of the Reissner–Nordström black hole. These results indicate that not only classical objects but also quantum wave functions are affected by gravity. This is obvious intuitively (classically), but not evident in quantum theory. We also studied the wave functions in the extremal and the overextremal cases.
@@
We then derived the Schrödinger equation for a molecule with a binary black hole as its nucleus. It is known that the Schrödinger equation for the hydrogen molecular ion $\mathrm{H}_{2}^{+}$ in flat spacetime can be solved exactly. We consider Majumdar–Papapetrou black holes as the nucleus. Since there is no force acting between these black holes, it is in some sense simpler than ordinary molecules. However, due to the presence of the curvature term, the spheroidal coordinates variables cannot be separated. Therefore, numerical analysis is generally required. We demonstrated that under specific conditions, the equation becomes separable and analytical solutions can be obtained.
@@
It would be helpful to know whether the black hole discussed in this paper is available in the current Universe or not. The Schwarzschild radius of a black hole of mass $M$ is given by $r_{0}=\frac{2G_{\mathrm{N}}M}{c^{2}}$, where $G_{\mathrm{N}}$ is Newton’s gravitational constant. On the other hand, the Bohr radius of the hydrogen atom is known as $a_{0}=\frac{4\pi\varepsilon_{0}\hbar^{2}}{m_{e}e^{2}}\sim 10^{-10}\,\mathrm{m}$ (where $m_{e}$ is the mass of the electron). The mass of the black holes of the size of the hydrogen atom is estimated as $M=\frac{c^{2}a_{0}}{2G_{\mathrm{N}}}\sim 6.7\times 10^{16}\,\mathrm{kg}$, which is of the same order as the small asteroid. Black holes of such a small mass are not created by the gravitational collapse of stars but may appear in the early stage of the Universe. The time $\Delta t$ that it takes for a black hole of mass $M\sim 10^{16}\,\mathrm{kg}$ to disappear through the Hawking radiation is estimated as $\Delta t\sim\frac{G_{\mathrm{N}}}{\hbar c^{3}}M^{3}\sim 10^{24}\,\mathrm{yr}$, which is greater than the age of the Universe. Therefore, it is possible that it remains in the current Universe as a primordial black hole. It cannot be excluded that such black holes trap electrons and form the “atoms” discussed in this paper.
@@
An interesting consideration related to what we discussed here is the discrepancy between the classical and quantum depictions of black hole atoms. In the purely classical picture, a test electron that is not supported by any nongravitational force and has insufficient angular momentum cannot remain at rest near a black hole. In such a situation, the electron cannot help crossing the horizon in a finite proper time along its timelike geodesic. In contrast, in the ordinary hydrogen atom in the flat space, quantum mechanics prevents the electron from collapsing into the nucleus and instead provides stationary bound states. Our results suggest an analogous mechanism for the “black hole atom,” namely, quantum effects can support bound states of the electron even in regions where classical motion admits no stable circular orbit so that the electron does not necessarily fall into the black hole. The “gravity versus quantum” effects is an interesting issue to study. It would be interesting to study this issue in more complex black holes, such as Kerr solutions and so on.
@@ Acknowledgments
The work of S. S. and K. S. is supported by Grant-in-Aid for Scientific Research, JSPS KAKENHI Grant Number JP25K07324. The work of K. S. is also supported by the Kitasato University Research Grant for Young Researchers.